the ratio of men to women in a certain factory is 3 to 4. there are 222 men. how many workers are there?

Answer by checkley79(3341)

3N=222
N=222/3
N=74 THIS IS THE MULTIPLIER FOR THE RATIO VALES.
4*74=296 THE NUMBER OF WOMEN.
222+296=518 TOTAL NUMBER OF WORKERS.
PROOF:
3/4=222/296
4*222=3*296
888=888